OpenStudy (anonymous):

Ive got the solution but i dont know the equation if how to solve this: My son is 1/3 my age now. Five years ago, he was 1/4 my age then. How old is my son?

OpenStudy (anonymous):

\[S = Son \\ G = Genuine \\ S = \frac{1}{3}G \\ \frac{1}{4}(G - 5) = S - 5\]

OpenStudy (anonymous):

OH thank you @failmathmajor !! I was really close to that though :L